Sudden and unexpected power cut at 10:02 German time, affecting all the target hall.
It has been caused by the emergency alarm being triggered. Reasons unclear. Electricians were informed.
CARME top gate valve closed.
10:33 power is back
Set point 120C
Scroll on, turbo on (980 Hz normal operation), valve open
Pressure at 1.34E-6 mbar
No obvious issues.
13:00
Power appears back for the whole ring.
Set point 128C
Turbo normal operation - no apparent adverse effects of shutdown
Pressure at ~3E-7 mbar
Pi can again be accessed by ssh. Vacuum plot is updating again. |